OpenStudy (anonymous):

5/11 divided by 13/22 is? answer has to be simplisted form

OpenStudy (anonymous):

invert the second fraction and multiply = 5/11 * 22/13 = 10/13

OpenStudy (anonymous):

\[5/11\div13/22\] \[5/11\times22/13\] 10/13

OpenStudy (anonymous):

appericate it

OpenStudy (anonymous):

medal plz click the good answer